OpenStudy (anonymous):

Find the exact value of the expression tan(sin^-1(4/9))

OpenStudy (anonymous):

\[\tan (\sin^{-1} 4/9)\]

OpenStudy (anonymous):

(4sqrt65)/65

OpenStudy (anonymous):

correct?

OpenStudy (lgbasallote):

was that what the calculator said?

OpenStudy (anonymous):

good point...but i'm not using a calc.

OpenStudy (lgbasallote):

you did that manually???? how could you!? :O

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

x^2+16=81 x^2=65 x=sqrt65 tan(theta)=opp/adj=4/sqrt65 =(4sqrt65)/65

OpenStudy (asnaseer):

your answer is correct

OpenStudy (anonymous):

ok, thanks

OpenStudy (asnaseer):

yw